A sample of concentrated HCL acid is 11.8M HCL and has a density of 1.190g/cm-cubed. Calculate: Mass fraction, Molality and Mole fraction.


1
Expert's answer
2021-09-19T00:01:58-0400

"\frac{11.8mol_{ HCl}}{L_{solution}}=\frac{mol_{HCL}}{Kg_{H_2O}}"


"\frac{1.190g_{solution}}{1ml_{solution}}×1000ml=1190g_{solution}"


"11.8mol\u00d736.5=430.7g"


Mass "=1190-430.7=759.3grams"


Mole fraction"=\frac{759.3}{1190}=0.638"